# Beyond Functional Correctness: Investigating Coding Style Inconsistencies in Large Language Models

### v1

@@ 4.2.3. Distribution of Coding Style Inconsistency Types
  • We then calculate the difference between the highest and lowest values of frequency of inconsistencies for each dimension. We sort the five dimensions from high to low according to the difference, and the result is: formatting (13.5%), fault tolerance (7.2%), statement/expression (6.2%), semantic (2.2%) and control flow (0.3%). This is because, although the training data of the models is generally similar, there are still some differences.
@@
RQ2 Summary: There are obvious coding style inconsistencies between human and all the studied Code LLMs. The top inconsistency type is API usage and top inconsistency dimensions are statements/expressions and formatting dimensions. While Code LLMs generally have similar coding styles, there are also noticeable differences in the formatting dimension.
@@ 4.3. RQ3: Coding Style Comparison
In addition to the analysis of coding style inconsistency between Code LLMs and human programmers, we further investigate which coding style is better. To this end, we annotate the code generated by Code LLMs by comparing it with the ground truth from three aspects: readability, conciseness, and robustness.
@@
  • Robustness: the ability of the code to handle corner cases and potential errors.
Based on the code samples generated by Code LLMs collected in RQ1, we compare them with the ground truth and score each of the three aspects according to the following criteria: model better (generated code is better than ground truth), tie (generated code is comparable to ground truth), and human better (the ground truth is better than the generated code). The annotation is conducted independently by two of the authors. Any conflicts are resolved through discussions to reach a consensus. Only valid code samples are considered for the annotation. Figure 11 show the proportion of code samples that received different scores (model better, tie, and human better) on the three aspects for each model.
@@
Overall, the code samples generated by the Code LLMs is comparable to that written by human programmers in terms of readability, conciseness, and robustness. On average, the code generated by the four models is comparable to or even superior to the code written by programmers in 86.2%, 79.9%, and 93.8% of cases in terms of readability, conciseness, and robustness, respectively. The following is a comparative analysis of the readability, conciseness, and robustness of the code samples generated by different Code LLMs. From the perspective of readability, the code samples generated by DeepSeekCoder-6.7B have the highest readability, while the code samples generated by CodeLlama-7B have the lowest readability. In terms of conciseness, the conciseness of code samples generated by CodeLlama-7B, StarCoder2-7B, and DeepSeekCoder-6.7B is comparable, while DeepSeekCoder-1.3B generates less concise code. Figure 10 presents an example that the conciseness of a code sample generated by DeepSeekCoder-1.3B is inferior to that of ground truth written by human programmers. Note that conciseness and readability are often trade-offs; in the example of Figure 11, DeepSeekCoder-1.3B makes the code more readable by splitting one statement into three statements. All four studied Code LLMs demonstrate relatively high robustness. This suggests that the models might have learned more robust coding styles from their training data, such as more rigorous input parameter checks, which human programmers might omit due to oversight or to avoid excessive complexity.
@@
RQ3 Summary: Overall, code generated by Code LLMs is comparable to or even better than human-written code in terms of readability, conciseness, and robustness. Among the studied models, DeepSeekCoder-6.7B produces the most readable code, while CodeLlama-7B and DeepSeekCoder-1.3B lags in readability and conciseness, respectively.
@@ 4.4. RQ4: Style Improvement by Prompting Techniques
In this RQ, we investigate whether prompting techniques can improve the coding style of Code LLMs. We conduct experiments with DeepSeekCoder-6.7B on 20 sampled Python tasks from CoderEval. We choose DeepSeekCoder-6.7B to conduct the experiment with type a because it achieves the best functional correctness in generating functions among the four models.These tasks are randomly selected from those that DeepSeekCoder-6.7B can complete, meaning DeepSeekCoder-6.7B can generate code samples that pass all corresponding test cases. We design four types of enhanced prompts for this study (refer to Figure  12), aiming to instruct the model to generate code with better coding style using explicit style guidelines. The design of these prompts investigates the impact of the placement and detail level of style guidelines. In prompt names, “-head” or “-end” specifies whether the style guidelines are placed before the function signature and docstring, similar to a directive, or appended at the end of the original docstring, simulating a normal docstring style. “-concise” and “-detailed” indicate the level of detail in the style guidelines. The detailed version includes three specific principles related to code readability, conciseness, and robustness, in addition to the concise information.
Among the selected tasks, DeepSeekCoder-6.7B generates 134 valid code samples using the basic prompt, i.e., the original function signature and docstring as input. Then, for each type of enhanced prompt, DeepSeekCoder-6.7B generates 10 code samples for the 20 selected tasks, resulting in 115, 137, 75, and 78 valid code samples for each of the four enhanced prompts, respectively. The accuracy for the four enhanced prompts is 57.5%, 68.5%, 37.5%, and 39.0%, respectively, compared to the 67.0% accuracy of the basic prompt. Except for prompt-head-detailed, the enhanced prompts result in lower accuracy compared to the basic prompt, suggesting that using more complex prompts may lead to a decrease in the functional correctness of the generated code.
@@
According to the scoring principles outlined in Section 4.3, we evaluated the code samples generated using the basic prompt and four enhanced prompts for readability, conciseness, and robustness. The results are depicted in Figure 13. Among the enhanced prompts, Prompt-head-concise, Prompt-end-concise, and Prompt-end-detailed slightly improve the readability of the code samples generated by DeepSeekCoder-6.7B. However, as shown in Figure 13(b), only Prompt-head-detailed enhances the conciseness of DeepSeekCoder-6.7B’s code samples. This is because there’s often a trade-off between readability and conciseness, where improving one may compromise the other. Additionally, as seen in Figure  13(c), all four enhanced prompts contribute to some extent to the improved robustness of DeepSeekCoder-6.7B’s code samples. In conclusion: (i) Incorporating style-guiding information into prompts may lead to decreased accuracy in generated code, as observed in our evaluation. (ii) Relying solely on prompt engineering may not fully resolve issues related to code style. Additional strategies or refinements may be necessary.
@@ 4.5. Case Studies
In the code samples we observed, we categorized and analyzed cases where the code samples exhibited inconsistent coding styles compared to the ground truth. We identified the following interesting scenarios.
Figure 14. Examples of (i) Using deprecated API; (ii) Unfamiliar with basic Python features; (iii) Rare use of advanced syntax features. On the left side of the image, (a), (c), and (e) are the ground truth for each example, while on the right side, (b), (d), and (f) are the code samples generated by Code LLMs.
Using deprecated APIs. In Figure 14 (b), the code sample generated by CodeLlama-7B uses the getchildren() method, which was deprecated in Python 3.2 and removed in Python 3.9. This might be due to CodeLlama-7B being trained on a corpus that includes Python code from different versions, leading to unawareness that certain APIs are outdated. Including deprecated APIs in code generated by large models is considered bad coding style, as this code will produce errors when run on newer Python versions.
Unfamiliar with basic Python features. Code LLMs might not be very familiar with some basic syntax features, which results in generating more complex code. For example, in Figure 14 (d), DeepSeekCoder-6.7B might not understand list slicing operations well, so it generated more complex code to avoid out-of-bounds indexing. Assuming the list has a length of 4, using list[3:5] in Python will not result in an error. Instead, it will return elements from index 3 to the end of the list. However, in the corresponding ground truth of the code sample (Figure 14 (c)), the code logic is clear and concise.
Rare use of advanced syntax features. Compared to code written by human programmers, code generated by Code LLMs often does not use advanced syntax features of the Python language, such as Pythonic idioms. As shown in Figure 14 (e), the ground truth uses list comprehension to build a list, while the code sample in Figure 14 (f), uses a more conventional method to build the list. It first constructs an empty list and then uses the append() method to add elements to the empty list. Compared to the ground truth, the simplicity of the code sample is inferior.
@@ 5. Threats to validity
Code LLM Utilization. Another potential threat is the utilization (e.g., source, parameter settings) of the Code LLMs used in our study. We carefully used the official release versions of each model to avoid any potential issues with unofficial or modified versions, followed the guidelines provided by the model developers to ensure proper implementation and usage, and conducted repeated tests to verify the performance and consistency of the models’ outputs. To ensure a fair comparison, we used the same prompt structure and generation parameters for each model, standardizing the experimental setup across different models.
Taxonomy Reliability and Completeness. The reliability and completeness of the inconsistency types identified pose another potential threat. We employed the open coding methodology to systematically identify and categorize inconsistency types, adhered to established open coding practices to ensure thoroughness and accuracy, and ensured that our terminology was stable by iteratively refining the inconsistency types until no new categories emerged. We involved multiple annotators to score these metrics, and they discussed their ratings to reach a consensus, reducing individual biases and ensuring more objective assessments. To further bolster the credibility of our findings, we have made all our data publicly available, allowing others to verify our results and methodology, thus enhancing the robustness of our conclusions.
@@ 6. CONCLUSION
Many studies have focused on improving the functional correctness of LLM-based code generation. However, the coding style of Code LLMs—an important aspect of code quality that extends beyond functional correctness—remains under-explored. To fill this gap, this paper makes the first attempt to investigate the coding style differences between LLMs and human developers through an empirical study. Specifically, we compare the code generation results of four mainstream Code LLMs with ground truth on the CoderEval benchmark.
We present a comprehensive taxonomy of coding style inconsistencies between Code LLMs and human developers, identifying 24 inconsistency types across five dimensions. Our analysis reveals clear coding style differences between the studied Code LLMs and human developers, particularly in statements/expressions and formatting, while showing similar coding styles among the studied Code LLMs. We further discuss potential causes of these style inconsistencies and explore ways to improve coding style discrepancies through prompt engineering, providing a foundation for future research in this area.
Acknowledgements.

### v2

@@ 4.2.3. Distribution of Coding Style Inconsistency Types
  • We then calculate the difference between the highest and lowest values of frequency of inconsistencies for each dimension. We sort the five dimensions from high to low according to the difference, and the result is: formatting (20.5%), semantic (7.3%), statement/expression (7.2%), fault tolerance (7.2%), and control flow (6.0%). This is because, although the training data of the models is generally similar, there are still some differences.
@@
RQ2 Summary: There are obvious coding style inconsistencies between human and all the studied LLMs. The top inconsistency type is API preference and top inconsistency dimensions are statements/expressions and formatting dimensions. While LLMs generally have similar coding styles, there are also noticeable differences in the formatting dimension.
@@ 4.3. RQ3: Coding Style Comparison
Figure 9. An example of how robustness relates to code style.
@@
In addition to analyzing coding style inconsistency between LLMs and human programmers, we further compare the code samples generated by LLMs and the ground truths in terms of three aspects: readability, conciseness, and robustness.
@@
  • Robustness: The ability of the code to handle corner cases and potential errors. To illustrate the relationship between robustness and code style, consider two programmers, A and B. Programmer B prefers to implement exception handling mechanisms in the code, while Programmer A tends to overlook this practice. In the Figure 10, func_A works correctly when the input is valid, but it will crash if the input is zero. In contrast, func_B includes an exception handling mechanism, ensuring the program does not crash and providing a meaningful error message when division by zero occurs. This demonstrates how Programmer B’s code is more robust due to different coding style.
We compare the code samples generated by LLMs from RQ1 with their corresponding ground truths and evaluate each of the three aspects. For each analysis, we determine whether the generated code is better than the ground truth (”model better”), whether they are comparable (”tie”), or whether the ground truth is better (”human better”). The scoring for the three aspects follows these principles: (i) Readability: If code sample A conveys the intent and logical structure more clearly than code sample B, it is superior in readability. (ii) Conciseness: If code sample A achieves the same functionality as code sample B while eliminating unnecessary redundancies, it is considered superior in terms of conciseness. This involves reducing the number of intermediate variables, optimizing loops, etc. (iii) Robustness: If code sample A effectively implements error handling and fault tolerance mechanisms, whereas code sample B lacks such mechanisms, the former is deemed superior in robustness. This includes the presence of input validation, exception handling, etc.
The annotation is conducted independently by two annotators. Both annotators have many years of programming experience and possess a certain level of understanding regarding research on code style. The two annotators independently score the code samples, resulting in two sets of annotations. Cohen’s Kappa (Cohen, 1960) is a statistical measure used to assess the consistency between two annotators for a classification task. We calculated the Cohen’s Kappa values for the two annotations on readability, conciseness, and robustness, which are 0.63, 0.79, and 0.68, respectively. However, the discrepancies that arose in the annotations were ultimately resolved through discussions to reach a consensus.
@@
Figure 11 show the proportion of code samples that received different scores (model better, tie and human better) on the three aspects for each model. Overall, the code samples generated by the LLMs is comparable to that written by human programmers in the three aspects. On average, the code generated by the five models is comparable to or even superior to the code written by programmers in 89.5%, 76.5%, and 94.2% of cases in terms of readability, conciseness, and robustness, respectively. The following is a comparative analysis of the readability, conciseness, and robustness of the code samples generated by the five LLMs. From the perspective of readability, the code samples generated by DeepSeekCoder-6.7B have the highest readability, while those generated by StarCoder2-7B have the lowest. In terms of conciseness, the conciseness of code samples generated by CodeLlama-7B, StarCoder2-7B, and DeepSeekCoder-6.7B is comparable, while GPT-4 generates less concise code. Figure 10 presents an example that the conciseness of a code sample generated by GPT-4 is inferior to that of ground truth written by human programmers. Note that conciseness and readability are often trade-offs. As shown in Figure 10, GPT-4 enhances readability by splitting one statement into two lines. All five studied LLMs demonstrate relatively high robustness. This suggests that the models might have learned more robust coding styles from their training data, such as more rigorous input parameter checks, which human programmers might omit due to oversight or to avoid excessive complexity.
@@
RQ3 Summary: Overall, the code samples generated by the five LLMs is comparable to human-written code in terms of readability, conciseness, and robustness. Among the studied models, DeepSeekCoder-6.7B produces the most readable code, while code samples generated by GPT-4 has the lowest simplicity but the highest robustness.
@@ 4.4. RQ4: Style Improvement by Prompting Techniques
In this RQ, we investigate whether prompting techniques can improve the coding style of LLMs. We conduct experiments with DeepSeekCoder-6.7B on 20 sampled Python tasks from CoderEval. These tasks are randomly selected from those that DeepSeekCoder-6.7B can complete, meaning DeepSeekCoder-6.7B can generate code samples that pass all corresponding test cases. We design four types of enhanced prompts for this study (refer to Figure  12), aiming to instruct the model to generate code with better coding style using explicit style guidelines. The design of these prompts investigates the impact of the placement and detail level of style guidelines. In prompt names, “-head” or “-end” specifies whether the style guidelines are placed before the function signature and docstring, similar to a directive, or appended at the end of the original docstring, simulating a normal docstring style. “-concise” and “-detailed” indicate the level of detail in the style guidelines. The detailed version includes three specific principles related to code readability, conciseness, and robustness, in addition to the concise information.
Among the selected tasks, DeepSeekCoder-6.7B generates 134 functionally correct code samples using the basic prompt, i.e., the original function signature and docstring as input. Then, for each type of enhanced prompt, DeepSeekCoder-6.7B generates 10 code samples for the 20 selected tasks, resulting in 115, 137, 75, and 78 functionally correct code samples for each of the four enhanced prompts, respectively. The accuracy for the four enhanced prompts is 57.5%, 68.5%, 37.5%, and 39.0%, respectively, compared to the 67.0% accuracy of the basic prompt. Except for prompt-head-detailed, the enhanced prompts result in lower accuracy compared to the basic prompt, suggesting that using more complex prompts may lead to a decrease in the functional correctness of the generated code.
@@
The two annotators mentioned in Section 4.3 referred to the scoring principles mentioned in Section 4.3 to score the code samples generated using the basic prompt and the four enhanced prompts for readability, conciseness, and robustness. The Cohen’s Kappa for the two sets of annotations on readability, conciseness, and robustness were 0.65, 0.81, and 0.74, respectively. Despite the discrepancies in the annotations, the annotators were eventually resolved through discussions, leading to a consensus. The G-test (Rohlf et al., 1981) is a statistical method used to assess the association between categorical variables. In this study, we applied the G-test to determine whether the probability distributions of scores for the basic prompt and the four enhanced prompts show significant differences across the three aspects of readability, conciseness, and robustness. The results indicated that the p-values for all three aspects were below 0.05, suggesting a statistically significant difference in the score distributions among the five prompts across these three aspects.
The results are depicted in Figure 13. Among the enhanced prompts, Prompt-head-concise, Prompt-end-concise, and Prompt-end-detailed slightly improve the readability of the code samples generated by DeepSeekCoder-6.7B. However, as shown in Figure 13(b), only Prompt-head-detailed enhances the conciseness of DeepSeekCoder-6.7B’s code samples. This is because there’s often a trade-off between readability and conciseness, where improving one may compromise the other. Additionally, as seen in Figure  13(c), all four enhanced prompts contribute to some extent to the improved robustness of DeepSeekCoder-6.7B’s code samples. In conclusion: (i) Incorporating style-guiding information into prompts may lead to decreased accuracy in generated code, as observed in our evaluation. (ii) Relying solely on prompt engineering may not fully resolve issues related to code style. Additional strategies or refinements may be necessary.
@@ 4.5. Case Studies
In the code samples we observed, we categorized and analyzed cases where the code samples exhibited inconsistent coding styles compared to the ground truth. We identified the following common scenarios.
Figure 14. Examples of (i) Poor formatting characteristics; (ii) Unfamiliar with basic Python features; (iii) Rare use of advanced syntax features. On the left side of the image, (a), (c), and (e) are the ground truth for each example, while on the right side, (b), (d), and (f) are the code samples generated by LLMs.
Poor formatting characteristics. LLMs may generate code with poor formatting characteristics, such as improper use of spaces. Figure 14 (b) is a function generated by GPT-4. In the expression “base_name.startswith(prefix+”-”)”, the “+” operator is not surrounded by spaces, which is not a good practice in the use of spaces. The code generated by the other four models also contains improper use of spaces, such as DeepSeekCoder-6.7B generating the code “yield response[start: end]”. However, the PEP-8 compliant version should be written as “yield response[start:end]”.
Unfamiliar with basic Python features. LLMs might not be very familiar with some basic syntax features, which results in generating more complex code. For example, in Figure 14 (d), DeepSeekCoder-6.7B might not understand list slicing operations well, so it generated more complex code to avoid out-of-bounds indexing. Assuming the list has a length of 4, using list[3:5] in Python will not result in an error. Instead, it will return elements from index 3 to the end of the list. However, in the corresponding ground truth of the code sample (Figure 14 (c)), the code logic is clear and concise.
Rare use of advanced syntax features. Compared to code written by human programmers, code generated by LLMs often does not use advanced syntax features of the Python language, such as Pythonic idioms. As shown in Figure 14 (e), the ground truth uses list comprehension to build a list, while the code sample in Figure 14 (f), uses a more conventional method to build the list. It first constructs an empty list and then uses the append() method to add elements to the empty list. Compared to the ground truth, the simplicity of the code sample is inferior.
@@ 5. Threats to Validity
LLM Utilization. Another potential threat is the utilization (e.g., source, parameter settings) of the LLMs used in our study. We carefully used the official release versions of each model to avoid any potential issues with unofficial or modified versions, followed the guidelines provided by the model developers to ensure proper implementation and usage, and conducted repeated tests to verify the performance and consistency of the models’ outputs. To ensure a fair comparison, we used the same prompt structure and generation parameters for each model, standardizing the experimental setup across different models.
Taxonomy Reliability and Completeness. The reliability and completeness of the inconsistency types identified pose another potential threat. We employed the open coding methodology to systematically identify and categorize inconsistency types, adhered to established open coding practices to ensure thoroughness and accuracy, and ensured that our taxonomy was stable by iteratively refining the inconsistency types until no new categories emerged. We involved multiple annotators to score these metrics, and they discussed their ratings to reach a consensus, reducing individual biases and ensuring more objective assessments. To further bolster the credibility of our findings, we have made all our data publicly available, allowing others to verify our results and methodology, thus enhancing the robustness of our conclusions.
@@ 6. Conclusion
Many studies have focused on improving the functional correctness of LLM-based code generation. However, the code style of LLMs—an important aspect of code quality that extends beyond functional correctness—remains under-explored. To fill this gap, this paper makes the first attempt to investigate the coding style differences between LLMs and human developers through an empirical study. Specifically, we compare the code generation results of five mainstream LLMs with ground truth on the CoderEval benchmark.
We present a comprehensive taxonomy of coding style inconsistencies between LLMs and human developers, identifying 24 inconsistency types across five dimensions. We adopt the definition of code style from previous work, but our taxonomy offers a more detailed classification of code style. Our analysis reveals clear coding style differences between the studied LLMs and human developers, particularly in statements/expressions and formatting, while showing similar coding styles among the studied LLMs. We further discuss potential causes of these style inconsistencies and explore ways to improve coding style discrepancies through prompt engineering, providing a foundation for future research in this area.
7. Data Availability
All the data used in this study is provided in the replication package (rep, 2024).
8. Acknowledgments
The work is supported by CCF-Huawei Populus Grove Fund CCFHuaweiSE202403, the Natural Science Foundation of Guangdong Province under Grant 2024A1515010255, and National Natural Science Foundation of China under Grant No. 62402113.